OpenStudy (anonymous):

*** WILL MEDAL *** A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6)?

OpenStudy (phi):

It looks like you keep subtracting eight: f(1)= 100 f(2) = f(1)-8 = 100-8 = 92 f(3) = f(2) -8 = 92 -8 = 84 and so on... the short way is to figure out f(6)= 100- 5*8